OpenStudy (anonymous):

A jet airliner moving initially at 300 mi/h due east enters a region where the wind is blowing at120mi/h in a direction 35.5 degrees north of east. What is the new velocity of the aircraft relative to the ground?

OpenStudy (zephyr141):



OpenStudy (zephyr141):

\[Resultant Vector= \sqrt{[120\cos(35.5)+300]^{2}+[120\sin(35.5)]^{2}}\]